6. Given h(x) x2-8x+ 21 a. What is the vertex of the parabola? b. What are the zeros of the function? Describe the type of zeros h(x) has. For example, does it have 2 real zeros, 2 complex zeros, 1 real zero, etc? Explain how you know this. c.
Answer #1

a)

h(x) =x²-8x+21

h(x) =(x² -8x +16) +5

h(x) =(x-4)² +5

Compare with standrad form h(x) =a(x-h)² +k to obatin vertex (h, k) as

(4,5)

b)

h(x) = 0

x²-8x+21 =0

2a 8±yB2-4.1.21 2.1

c)

h(x) has two complex zeroes, as both contain imaginary term i.